OpenStudy (anonymous):

What comes next in the pattern 1.5, 2.75, 4,...

OpenStudy (anonymous):

5.5

OpenStudy (anonymous):

- how did you get that ?

OpenStudy (anonymous):

It goes, 1.5, 2.75, and 4, next would be 5.5 if it would go on further, it would be, 1.5 2.75 4 5.5 6.75 8 and so on.

OpenStudy (anonymous):

Thankss . I need help with 2 more lol .

OpenStudy (anonymous):

No problem! :) What are the 2 more?

OpenStudy (anonymous):

& i have to explain how i got that ^😐

OpenStudy (anonymous):

Is it a multiple choice answer?

OpenStudy (anonymous):

- no lol . It just say " description of how you figured out the pattern"

OpenStudy (anonymous):

Would my answer be a good one?

OpenStudy (anonymous):

You keep adding 1.25

OpenStudy (anonymous):

- thanks @ tessa

OpenStudy (anonymous):

- what would come next after 1/8, 1/4, 1/2'... & is it increasing or decreasing ?

OpenStudy (anonymous):

The first one is 5.25 not 5.50 because you are increasing by 1.25

OpenStudy (anonymous):

-thankss

OpenStudy (anonymous):

1/8=12.5% 1/4=25% 1/2 =50%. Hopefully this helps